OpenStudy (anonymous):

What is the solution set for |–4x – 4| – 28 = –4?

OpenStudy (mertsj):

Add 28 to both sides. What do you get?

OpenStudy (anonymous):

24

OpenStudy (mertsj):

Good.

OpenStudy (anonymous):

I feel like I should let you explain step by step instead of blurting out the answer....

OpenStudy (mertsj):

|4x-4|=24 means that 4x-4 is exactly 24 units from 0.

OpenStudy (mertsj):

That means that 4x-4 could be 24 or 4x-4 could be -24

OpenStudy (mertsj):

\[4x-4=24 or 4x-4=-24\]

OpenStudy (mertsj):

So you solve both of those equations. Can you do that?

OpenStudy (anonymous):

these are the awser that they give me x = –7 and x = 5 x = –7 and x = –5 x = 7 and x = 5 x = 7 and x = –5

OpenStudy (mertsj):

Which one is it?

OpenStudy (mertsj):

what do you get when you solve 4x-4=24?

OpenStudy (anonymous):

i am not sure

OpenStudy (mertsj):

Add 4 to both sides. What do you get?

OpenStudy (mertsj):